# Out-of-equlibrium inference of stochastic model parameters through population data from generic consumer-resource dynamics

### v1

@@ Abstract
Consumer-resource dynamics is central in determining biomass transport across ecosystems. The assumptions of mass action, chemostatic conditions and stationarity in stochastic feeding dynamics lead to Holling type II functional responses, whose use is widespread in macroscopic models of population dynamics. However, to be useful for parameter inference, stochastic population models need to be identifiable, this meaning that model parameters can be uniquely inferred from a large number of model observations. In this contribution we study parameter identifiability in a multi-resource consumer-resource model, for which we can obtain the steady-state and out-of-equilibrium probability distributions of predator’s abundances by analytically solving the master equation. Based on these analytical results, we can conduct in silico experiments by tracking the abundance of consumers that are either searching for or handling prey, data then used for maximum likelihood parameter estimation. We show that, when model observations are recorded out of equilibrium, feeding parameters are truly identifiable, whereas if sampling is done at stationarity, only ratios of rates can be inferred from data. We discuss the implications of our results when inferring parameters of general dynamical models.
@@
stochastic consumer-resource models, master equation, out-of-equilibrium distribution, parameter inference, multi-resource Holling type II feeding dynamics
I Introduction
Ecological communities are systems formed by individuals of different species interacting with each other, through a variety of mechanisms [1], as well as with the external environment [2]. The mathematical description of these systems amounts to provide quantitative predictions of species abundances as functions of time, together with their spatial variability at different scales. The gradual acknowledgement that species interactions in ecological communities can be well described mathematically [3] has resulted in making Ecology an ever-increasingly quantitative science. Among all the interactions that shape the complexity of ecological communities, the consumer-resource interaction is of particular relevance as it structures the way biomass is transported across the ecosystem [4, 5, 6]. Determining who eats whom and how allows ecologists to describe ecological communities as complex networks (food webs) where the flow of biomass between different species can be quantified [7, 8].
@@ I Introduction
In classical approaches, biomass flow is quantified by inter-dependent changes in species abundances, which are usually measured by the density of individuals (or their biomass) of each species in a given region [9]. This quantification has relied mainly on differential equations describing the spatial and temporal variability of species abundances. Consumer-resource interactions are included in these equations via mathematical functions called functional responses, which describe the mechanisms associated to feeding interactions [10, 11]. Functional responses are defined as the per capita, average feeding rate of consumers on their resources [12]. Such functions have been derived both from conceptual considerations about the feeding dynamics (ranging from simple predator-prey equations to more complex food-web models [13, 11]), but also from feeding experiments [14, 15, 11] where a consumer is observed feeding on a resource population. Functional responses have been generalized to multi-species systems [16, 17] and to include a variety of foraging mechanisms.
Holling’s seminal papers [18, 19] show how the behavior of individual consumers gives rise to an average feeding rate over the whole consumer population, which then can be used in population-level models. In line with this, the way functional responses are inferred from feeding experiments [14, 20] makes their use in population dynamics more consistent [21]. The combined effect of elemental processes at a lower level of aggregation (e.g., the individual level) can be averaged to yield a macroscopic dynamics at a higher level of aggregation (e.g., the population or the community level). This methodology, which is widespread in Statistical Mechanics [22], has been extended to population and community ecology [23]. Individuals are regarded as interacting particles, whose interaction is described in terms of elemental processes as in chemical reaction kinetics [24]. The macroscopic dynamics (i.e., the set of differential equations that describe the temporal variation of species abundances) then emerges as the aggregated effect of the stochastic processes operating at the individual level.
@@
Elemental processes in consumer-resource models are driven by feeding parameters. In the simplest scenario one can imagine, the feeding process involves two mechanisms. On the one hand, the attack of searching predators to available prey, to form compounds in which the predator is handling the prey and not searching for other resources. On the other hand, the predator-prey pair takes some (handling) time, which represents the process of handling and digesting the prey item by the individual consumer. Both attack and relaxation processes occur at certain rates which are not known a priori. These rates could be inferred from population abundances of the different types involved in the feeding process, given some hypotheses about how these abundances vary according to the two elemental processes. In particular, the distributions of the population numbers of each behavioral type (free and handling consumer numbers, in the example above) can be used to infer feeding parameters, provided that the statistical model is fitted to empirical abundance data. In this contribution, we focus on the problem of model parameters inference using stochastic population models and in silico abundance data.
Stochastic models are to be designed so that feeding parameters can be uniquely estimated from experimental abundance data. However, precise statistical inference of model parameters cannot be done for every statistical model: a model is known as identifiable if it is theoretically possible to infer true parameter values after retrieving an infinite number of observations from the model. Mathematically, this condition amounts to saying that different parameter sets lead to strictly different distributions of observable quantities pertaining to the sample space. More precisely, let $\mathcal{G}=(\mathcal{S},\mathcal{P})$ be a statistical model with sample space $\mathcal{S}$, onto which observable variables $x$ take values, and a set of probability distributions $P_{\theta}(x)$ whose parameter set $\theta$ is a member of the parameter space $\Theta$:
$\mathcal{P}=\{P_{\theta}|\theta\in\Theta\}.$ (1)
Model $\mathcal{G}$ is identifiable if the mapping $\theta\mapsto P_{\theta}$, which assigns a probability distribution to a set of parameters, is one-to-one [25]. As said, this means that distinct parameter sets should correspond to distinct probability distributions, i.e.: $P_{\theta_{1}}(x)=P_{\theta_{2}}(x)\Leftrightarrow\theta_{1}=\theta_{2}$ almost surely.
For an identifiable model, the map $\theta\mapsto P_{\theta}$ can be inverted to learn the true parameter set if the model can be observed an infinite number of times. This is true because of the strong law of large numbers, by which probabilities can be estimated by averaging frequencies over a large number of observations [26]. Thus, with an infinite number of observations, finding the true probability distribution $P_{\theta_{0}}$ is feasible. Because the map $\theta\mapsto P_{\theta}$ is invertible, the true parameter set $\theta_{0}$ which generated the distribution $P_{\theta_{0}}$ can be unambiguosly extracted from the distribution.
@@
Holling type II functional response can be obtained only when mass action (individuals meet at random), chemostatic conditions (the total number of consumers and the total number of resources are kept constant) and stationarity are assumed [27, 28]. In this contribution, we focus on the problem of model identifiability using a multi-resource, Holling type II feeding dynamics. For that purpose, we consider a single predator population, in which individuals are able to search and handle $S$ different resource species, leading to $S+1$ different behavioral types, namely: $S$ types associated to predators handling each resource species, and free predators that are not handling any resource. We assume mass action and chemostatic conditions to derive the joint probability distribution of the different behavioral types, at two regimes: in the transient dynamics (out of equilibrium) and in the long run (at stationarity). We show analytically that, at equilibrium, the model is not identifiable because only the ratio between relaxation and attack rates can be extracted from model fitting to data. In this sense, the model at equilibrium is set identifiable, this meaning that parameters can be precisely determined only on a subset of the parameter space. We also show that true feeding parameters cannot be properly estimated via a standard fitting procedure such as maximum likelihood (ML) parameter estimation using synthetic (model-simulated) data. This in silico experiment, which mimics an empirical setup to estimate feeding parameters that assumes the system has reached stationarity, only allows for precise estimation of the quotient of rates at equilibrium.
@@
However, if model observations are taken out of equilibrium at different sampling times, feeding parameters can be uniquely estimated via ML. In this case, the likelihood function associated to independent observations of the system’s state exhibits a maximum at precise parameter values, making the model completely identifiable. This allows for proper parameter inference from data, providing a practical solution for the absence of idenfiability at equilibrium. In this line, to avoid potential ambiguities in dynamical model calibration, we stress that parameter inference should be done out of equilibrium. We also show that, for the multi-resource feeding dynamics studied, model identifiability is lost if sampling is done for independent dynamical observations taken at the exact same time (either close to equilibrium or not). In actual experiments, this is not a shortcoming for inference. The degeneracy of optima in the likelihood function for equal sampling times explains why the model is not identifiable at equilibrium.
In order to carry out the analysis, we solve the stochastic model at equilibrium by computing the joint probability distribution of behavioral types, which turns out to be a multinomial distribution. Moreover, we solve the master equation to obtain the temporal dependence of the joint probability distribution in the case of equal relaxation rates, which leads also to a multinomial with probability parameters depending on time. Thanks to these calculations, we can conduct parameter inference, both at equilibrium and out-of-equilibrium, using synthetic datasets based on model observations. Moreover, the model analysis can be extended to different scenarios in biophysical systems other than pure consumer-resource dynamics. For example, the multi-resource Holling type II dynamics solved here can be applied to model the flux of molecules occurring during transport across the cell membrane (Figure 1). Generic membrane proteins acting as transporters are the counterpart of consumers (which, in general, will have a different affinity for each type of molecule). Transporters can be in a “handling” state, containing inside the different molecules (i.e., different resources, such as nutrients), which are released into the cytoplasm after certain relaxation time has elapsed, moment after which the transporter comes back to a free state. Our mathematical derivations can be used to estimate the distribution of free and handling transporters as function of time, given that chemostatic conditions (constant density of each molecule type) apply.
Figure 1: Modeling cellular nutrient transport as an ecological consumer-resource model. The multi-resource consumer-resource model we analyze in this contribution can be applied to other biophysical systems, for example, to model nutrient transport in the cell membrane. Proteins acting as transporters (orange rectangles) are regarded as consumers, which can be free (with no nutrient inside; for example, the second and fifth ones) or “handling” inside different types of nutrients, which are regarded as different resources. Two processes are involved: the formation of the compound transporter-nutrient, which occurs at a certain “attack” rate, and the breakage of the compound leading to nutrient release into the cytoplasm, which takes place at a given relaxation rate. Both rates can be nutrient-dependent.
@@
The paper is organized as follows. The analytical solution of the stochastic model is explained in detail in Section II. Section III is devoted to analyze several in silico experiments designed to illustrate model identifiability, showing that feeding parameters can be only fully identifiable if the model is sampled out of equilibrium at different sampling times. At equilibrium, the model allows for precise determination of dimensionless parameter ratios. Finally, we conclude discussing the implications and potential extensions of our work.
II Multi-resource Holling Type II feeding dynamics
We start by defining the consumer-resource dynamics. Assume that an individual consumer $X_{A}$ can feed on a range of $S$ alternative resource species, $X_{R_{i}}$, with $i=1,\dots,S$. We consider two processes: first, the attack of resource $X_{R_{i}}$ by the consumer $X_{A}$, which results in the formation of a compound $X_{AR_{i}}$ formed by the consumer handling the resource. Second, handling consumers do not search for prey until the compound $X_{AR_{i}}$ “relaxes” because the predator stops handling the prey, then becoming a consumer searching for prey. The stochastic dynamics can be summarized by $2\,S$ individual feeding reactions:

### v2

# Out-of-equlibrium inference of feeding rates through population data from generic consumer-resource stochastic dynamics

@@ Abstract
Statistical models are often structurally unidentifiable, because different set of parameters can lead to equal model outcomes. To be useful for prediction and parameter inference from data, stochastic population models need to be identifiable, this meaning that model parameters can be uniquely inferred from a large number of model observations. In particular, precise estimation of feeding rates in consumer-resource dynamics is crucial, because consumer-resource processes are central in determining biomass transport across ecosystems. Model parameters are usually estimated at stationarity, because in that case model analyses are often easier. In this contribution we analyze the problem of parameter redundancy in a multi-resource consumer-resource model, showing that model indentifiability depends on whether the dynamics have reached stationarity or not. To be precise, we: (i) Calculate the steady-state and out-of-equilibrium probability distributions of predator’s abundances analytically using generating functions, which allow us to unveil parameter redundancy and carry out proper maximum likelihood estimation. (ii) Conduct in silico experiments by tracking the abundance of consumers that are either searching for or handling prey, data then used for maximum likelihood parameter estimation. (iii) Show that, when model observations are recorded out of equilibrium, feeding parameters are truly identifiable, whereas if sampling is done at stationarity, only ratios of rates can be inferred from data (i.e., parameters are redundant). We discuss the implications of our results when inferring parameters of general dynamical models.
keywords:
@@
Stochastic consumer-resource models , Master equation , Model identifiability , Parameter inference , Generating function , Multi-resource Holling type II and III feeding dynamics
journal: Applied Mathematics and Computation
1 Introduction
@@
The consumer-resource interaction is a key driver of the transport of biomass across ecological communities (Belgrano, 2005; Dunne, 2006; Getz, 2011; Perkins et al., 2022; Valdovinos et al., 2022). In classical approaches, biomass flow is quantified by inter-dependent changes in species abundances, which are usually measured by the density of individuals of each species (or their biomass) in a given region (Bersier et al., 2002). This quantification has relied mainly on differential equations describing the spatial and temporal variability of species abundances. Consumer-resource interactions are included in these equations via mathematical functions called functional responses, which encapsulate and simplify the complexities of feeding interactions, providing a coarse-grained representation of the underlying mechanisms (Lafferty et al., 2015; Abrams, 2022). Functional responses are defined as the per capita, average feeding rate of consumers on their resources (Berryman, 1992). Such functions have been derived from conceptual considerations about feeding dynamics (ranging from simple predator-prey equations to more complex food-web models (McPeek, 2019; Abrams, 2022)), but also from feeding experiments (Rosenbaum and Rall, 2018; Barbier et al., 2021; Abrams, 2022), where a consumer is observed feeding on a resource population. Holling’s seminal papers (Holling, 1959a, b) show how the behavior of individual consumers gives rise to an average feeding rate over the whole consumer population, which then can be used in population-level models. As shown in Palamara et al. (2021, 2022), Holling’s type II functional response can be rigorously derived from processes at the individual level when mass action (individuals meet at random), chemostatic conditions (the total number of consumers and the total number of resources are kept constant) and stationarity are assumed. Functional responses have been generalized to multi-species systems (Koen-Alonso, 2007; Robertson et al., 2022) and to include a variety of foraging mechanisms.
@@
Elemental processes at the lower level of aggregation (i.e., individual level) in consumer-resource models are driven by feeding parameters. In the simplest scenario one can imagine, the feeding process involves two predator behavioral stages: the search for and the handling of preys. The search or hunting behavior of the available prey ends with a predator attack, moment at which the predator starts handling the prey and stops searching for other preys. This predator-prey complex is usually irreversible (unless the prey escapes) and relaxes into a free predator again ready to search for more prey. This relaxation process takes some (handling) time, during which the individual predator kills, handles, and diggest the prey. Both attack and relaxation processes occur at certain rates that are not known a priori. These rates could be inferred from population abundances of the different types involved in the feeding process, given some hypotheses about how these abundances vary according to the two elemental processes. In particular, the distributions of the population numbers of each behavioral type (free and handling consumer numbers, in the example above) can be used to infer feeding parameters, provided that the statistical model is fitted to empirical abundance data. In this contribution, we focus on the problem of parameter inference using stochastic population models and in silico (simulated) abundance data.
@@
Stochastic models are to be designed so that model parameters can be uniquely determined from a large number of model realizations. In that case, the model is known as structurally identifiable, because model parameters can be inferred without redundancy from data, regardless the quantity or quality of the data (Bellman and Åström, 1970). In particular, dynamic consumer-resource models should allow per-capita feeding rates (associated to elemental feeding processes) to be estimated unambiguously from experimental population abundance data. The problem of structural identifiability has been thoroughly analyzed for deterministic dynamics in the context of general biological modeling approaches (see Villaverde et al. (2016) and references therein). In this contribution, we focus on the problem of structural identifiability in stochastic dynamical models. Our aim is to unveil scenarios for feeding rate redundancy using a multi-resource, Holling type II feeding dynamics (Holling, 1959a, b). For that purpose, we consider a single predator population, in which individuals are able to search for and handle $S$ different resource species, leading to $S+1$ different behavioral types, namely: $S$ types associated to predators handling each resource species, and free predators that are not handling any resource. We assume mass action and chemostatic conditions to derive the joint probability distribution of the different behavioral types, at two regimes: in the transient dynamics (out of equilibrium) and in the long run (once the steady-state is reached). We show analytically that, at equilibrium, the model is not identifiable because only the ratio between relaxation and attack rates can be extracted from fitting model to data. In this sense, the model at equilibrium is set identifiable, this meaning that parameters can be precisely determined only on a subset of the parameter space. We also show that true feeding rates cannot be properly estimated via a standard fitting procedure such as maximum likelihood (ML) parameter estimation using synthetic (model-simulated) data at stationarity. This in silico experiment, which mimics an empirical setup to estimate feeding parameters that assumes the system has reached stationarity, only allows for precise estimation of quotients of rates at equilibrium.
@@
However, if model observations are taken out of equilibrium at different sampling times, feeding parameters can be uniquely estimated via ML. In this case, the likelihood function associated to independent observations of the system’s state exhibits a maximum at precise parameter values, making the model completely identifiable. This allows for proper parameter inference from data, providing a practical solution for parameter redundancy at equilibrium. In this line, to avoid potential ambiguities in model calibration from time series data, we stress that parameter inference should be done out of equilibrium. We also show that, for the multi-resource feeding dynamics studied, parameter redundancy arises if sampling is done for independent dynamical observations taken at the exact same time (either close to equilibrium or not). In actual experiments, this is not a shortcoming for inference.
@@
In order to carry out the analysis, we solve the stochastic model at equilibrium by computing the joint probability distribution of behavioral types, which turns out to be a multinomial distribution. In addition, we solve the master equation to obtain the temporal dependence of the joint probability distribution in the case of equal relaxation rates, which leads also to a multinomial with probability parameters depending on time. Thanks to these calculations, we can conduct parameter inference, both at equilibrium and out-of-equilibrium, using synthetic datasets based on model observations. The paper is organized as follows. Section 2 summarizes the background on parameter redundancy and shows that the explicit computation of the generating function can be used to assess parameter redundancy. The analytical solution of the stochastic model is explained in detail in Section 3. Section 4 is devoted to analyze several in silico experiments designed to illustrate parameter identification failure, showing that feeding parameters can be only fully identifiable if the model is sampled out of equilibrium at different sampling times. At equilibrium, the model allows for the precise determination of dimensionless parameter ratios. Finally, we conclude discussing the implications and potential extensions of our work.
2 Stochastic model identifiability via generating functions
Precise inference of model parameters cannot be done for every statistical model: a model is defined as identifiable if it is theoretically possible to infer true parameter values after retrieving an infinite number of observations from the model. Mathematically, this condition amounts to saying that different parameter sets lead to strictly different distributions of observable quantities pertaining to the sample space. More precisely, let $\mathcal{G}=(\mathcal{S},\mathcal{P})$ be a statistical model with sample space $\mathcal{S}$, onto which a vector of observable variables $\bm{x}$ takes values, and a set $\mathcal{P}$ of probability distributions $P_{\bm{\theta}}(\bm{x})$ whose parameter vector $\bm{\theta}$ is a member of the parameter space $\Theta$:
$\mathcal{P}=\{P_{\bm{\theta}}|\bm{\theta}\in\Theta\}.$ (1)
Model $\mathcal{G}$ is identifiable if the mapping $\bm{\theta}\mapsto P_{\bm{\theta}}$, which assigns a probability distribution to a set of parameters, is one-to-one (Lehmann and Casella, 2006). As said, this means that distinct parameter sets should correspond to distinct probability distributions, i.e.: $P_{\bm{\theta}_{1}}(\bm{x})=P_{\bm{\theta}_{2}}(\bm{x})\Leftrightarrow\bm{\theta}_{1}=\bm{\theta}_{2}$ almost surely for all $\bm{x}$.
For an identifiable model, the map $\bm{\theta}\mapsto P_{\bm{\theta}}$ can be inverted to learn the true parameter set if the model can be observed an infinite number of times. This is true because of the strong law of large numbers, by which probabilities can be estimated by averaging frequencies over a large number of observations (Loève and Loève, 1977). Thus, with an infinite number of observations, finding the true probability distribution $P_{\bm{\theta}_{0}}$ is feasible. Because the map $\bm{\theta}\mapsto P_{\bm{\theta}}$ is invertible, the true parameter set $\bm{\theta}_{0}$ which generated the distribution $P_{\bm{\theta}_{0}}$ can be unambiguosly extracted from the distribution.
The easiest way to check for parameter redundancy is by finding an exhaustive summary (Cole, 2020), which is a vector $\bm{\kappa}(\bm{\theta})$ of model parameter combinations that uniquely determine the statistical model $\mathcal{G}$, so that $\bm{\kappa}(\bm{\theta}_{1})=\bm{\kappa}(\bm{\theta}_{2})\Leftrightarrow P_{\bm{\theta}_{1}}(\bm{x})=P_{\bm{\theta}_{2}}(\bm{x})$ for all $\bm{\theta}_{1}$, $\bm{\theta}_{2}\in\Theta$. In general, any model will have many suitable exhaustive summaries. Knowledge of values of the exhaustive summary fully determines model outcomes. If the parameter vector $\bm{\theta}$ can be determined uniquely from an exhaustive summary $\bm{\kappa}(\bm{\theta})$, then the model lacks of parameter redundancy. See Castro and De Boer (2020) for more advanced ways of testing structural identifiability.
Stochastic model identifiability can be assessed if one can compute the distribution $P_{\bm{n}}(t)$ of observing state $\bm{n}$ at time $t$, which will depend on model parameters. For that purpose, the master (forward Kolmogorov) equation has to be solved (Gillespie, 1991), and from $P_{\bm{n}}(t)$ the exhaustive summary could be eventually extracted. Equivalently, one can obtain the multivariate $z$-transform of $P_{\bm{n}}(t)$, known as generating function, defined as
where the vector $\bm{z}=(z_{1},\dots,z_{S})$ is conjugate to the state vector $\bm{n}=(n_{1},\dots,n_{S})$. This vector $\bm{z}$ can take values over $\mathbb{R}^{S}$. A series expansion will eventually lead to the probabilities $P_{\bm{n}}(t)$. Having $G(\bm{z},t)$ is equivalent to knowing $P_{\bm{n}}(t)$, because the generating function allows to compute every moment of the probability distribution, thus uniquely determining it. Therefore, a set of parameter combinations that uniquely determine the generating function can be used as an exhaustive summary to assess parameter redundancy.
We illustrate this idea using a well-known birth-death-immigration stochastic dynamics for a single population (Kendall, 1949). The elemental processes are defined by the following reaction scheme:
@@ 2 Stochastic model identifiability via generating functions
[…]
The first process stands for the arrival of a new individual $X$ by immigration at rate $\mu$, irrespectively of the number of individuals present in the system. The second one refers to reproduction at a per-capita rate $\beta$, this meaning that if there are $n$ individuals, they will reproduce with probability $\beta n$ per unit time. Finally, individual death occurs at a per-capita rate $\delta$. Since birth and death rates are linear on population numbers, the master equation can be solved analytically in terms of the generating function
$G(z,t)=\sum_{n\geq 0}P_{n}(t)z^{n},$ (6)
which results in (Kendall, 1949)
$G(z,t)=\frac{\left(\frac{\beta}{\delta}-1\right)^{\frac{\mu}{\beta}}\left[e^{(\beta-\delta)t}(1-z)+\frac{\beta}{\delta}z-1\right]^{n_{0}}}{\left[\frac{\beta}{\delta}e^{(\beta-\delta)t}(1-z)+\frac{\beta}{\delta}z-1\right]^{n_{0}+\frac{\mu}{\beta}}}$ (7)
for an initial condition assuming $n(0)=n_{0}$ individuals at time $t=0$. We focus on the problem of estimating elemental process rates ($\beta$, $\delta$, and $\mu$) and assume that the initial condition $n_{0}$ is known in advance, so we do not need to infer it from data. The generating function —and, henceforth, the probability $P_{n}(t)$— is uniquely determined by specifying the parameter combinations $\mu/\beta$, $\beta-\delta$, and $\beta/\delta$. Thus, an exhaustive summary is $\bm{\kappa}(\bm{\theta})=(\mu/\beta,\beta-\delta,\beta/\delta)$. If vector $\bm{\kappa}(\bm{\theta})$ is known, then the parameter vector $\bm{\theta}=(\beta,\delta,\mu)$ can be uniquely determined and the model is identifiable. Therefore, the whole parameter set can be inferred from model observations via maximum likelihood parameter estimation (MLE) techniques, for example.
On the other hand, let $\beta<\delta$ and assume that model observations are conducted at stationarity ($t\to\infty$). The generating function reduces to
$G(z,\infty)=\left(1-\frac{\beta}{\delta}\right)^{\frac{\mu}{\beta}}\left(1-\frac{\beta}{\delta}z\right)^{-\frac{\mu}{\beta}}.$ (8)
Besides the trivial issue that $n_{0}$ cannot be inferred from steady-state observations in models with a single attractor like this one, there is a major issue in determining elemental process rates. According to (8), an exhaustive summary that fully determines steady-state model realizations is $\bm{\kappa}(\bm{\theta})=(\mu/\beta,\beta/\delta)$ for $\bm{\theta}=(\beta,\delta,\mu)$. Therefore, only the ratios $\mu/\beta$ and $\beta/\delta$ are truly identifiable, and the model remains parameter redundant (set identifiable) in this limit.
This simple example illustrates a problem that we expect to occur frequently, in stochastic population models, when it comes to estimating individual process rates. First, the master equation cannot be solved for general, multi-species birth-death processes when transition probability rates depend non-linearly on population numbers. Second, among the solvable models, it is much more frequent to find steady-state solutions of the master equation (thanks to detailed-balance conditions, for example) rather than the general temporal dependence of $P_{\bm{n}}(t)$. Therefore, having computed the steady-state solution $P_{\bm{n}}$ analytically, one could naively formulate a MLE method to estimate model rates by sampling the model at stationarity. As the example shows, this can lead to infinite possible estimates. Although model samples are generated at precise values of elemental process rates, they cannot be inferred from stationary distributions. The issue precisely stems by sampling the model after the steady state has been reached.
We conjecture this issue to be general (at least for models defined by elemental process rates with a polynomial dependence on population abundances), because at equilibrium there is no well-defined temporal scale, and the generating function could depend only on ratios of rates. Therefore, parameters should be estimated out of equilibrium. This has important implications for population models that are not analytically tractable, for which numerical log-likelihood profiles or simulation are to be applied. Moreover, if the MLE procedure is based on abundance data, one should be aware of whether data were collected at stationarity or not to be truly confident of ML estimates. In the rest of the paper, we demonstrate this pathology of MLE at equilibrium using a solvable multi-species consumer-resource model and show that the issue is resolved when data for MLE is collected out of equilibrium. In the discussion section, we highlight that natural extensions of this model would also be parameter redundant at equilibrium.
3 Multi-resource Holling Type II feeding dynamics
We start by defining a model for consumer-resource dynamics. Assume that an individual consumer $X_{A}$ can feed on a range of $S$ alternative resource species, $X_{R_{i}}$, with $i=1,\dots,S$. We consider two processes: first, the attack of resource $X_{R_{i}}$ by the consumer $X_{A}$, which results in the formation of a compound $X_{AR_{i}}$ formed by the consumer handling the resource. Second, handling consumers do not search for prey until the compound $X_{AR_{i}}$ “relaxes” because the predator stops handling the prey, then becoming a consumer searching for prey. The stochastic dynamics can be summarized by $2\,S$ individual feeding reactions:
[…]